# On the arithmetic of polynomial ideals

### v1

@@ Abstract.
This research was funded in whole, or in part, by the Austrian Science Fund (FWF) [10.55776/DOC183 and 10.55776/ PAT9756623]. L. C. also acknowledges the European Union’s Horizon 2020 program (Marie Sklodowska-Curie grant 101021791) for the financial support on the early stages of the project. L. C. is a member of the National Group for Algebraic and Geometric Structures and their Applications (GNSAGA), a department of the Italian Mathematics Research Institute (INdAM)
@@ 1. Introduction
Our work contributes to the extensive literature on monoids of ideals: we refer to [29] and the monograph [17] for an overview, and to [33, 1, 30, 31] for more recent develpoments. A common approach in the literature is to focus on specific classes of ideals within a given ring, or semigroup ideals within a given monoid. Among these, divisorial ideals are perhaps the most prominent example. By suitably choosing a ring (or monoid) and a class of its ideals, one can obtain a fairly well-behaved monoid of ideals, and in some cases one is able to completely determine its structure. As an example, we recall two well-known results: a domain $R$ is Dedekind if and only if all its nonzero ideals are invertible, and this happens only if the monoid of all nonzero ideals is factorial, and a domain $R$ is Krull if and only if the monoid of its invertible divisorial ideals is factorial.
@@
In the paper [21], a first step was made in the study of the arithmetic properties of ${\mathcal{I}}(R)$ when $R$ a multivariate polynomial ring. In particular, the results collected in [21, Theorem 5.1] reveal a noticeable similarity between the arithmetic of $\mathcal{I}(R)$ and that of Krull monoids with infinite class group and prime divisors in each class, despite ${\mathcal{I}}(R)$ not even being transfer Krull. More specifically, ${\mathcal{I}}(R)$ is fully elastic and each of its unions of sets of lengths is the whole set $\mathbb{N}_{\geq 2}$.
Building on these initial results, we refine the demonstrative techniques and adapt them to broader contexts. Following a review of the necessary preliminaries in Section 2, we extend the methods introduced in [21] to find several new families of atoms in ${\mathcal{I}}(R)$. In particular, we highlight in in Section 3.1 an interesting class built from sum-free subsets of $\mathbb{N}$, which includes some previously known examples.
Motivated by the problem of computing the length sets of certain ideals (see Theorem 3.7), in Section 4 we focus on the submonoid ${\mathcal{M}{\rm on}}(R)\subseteq{\mathcal{I}}(R)$ of monomial ideals of $R$ (i.e., ideals generated by monomials). We first provide some arithmetic results in the spirit of [21, Theorem 5.1], and then we determine several classes of atoms in this monoid, allowing us to compute the set of lengths of the families of ideals introduced in Theorem 3.7. The main technical ingredient which makes the study of factorizations in ${\mathcal{M}{\rm on}}(R)$ easier than in ${\mathcal{I}}(R)$ is Lemma 4.1 (see also Remark 4.4) greatly constricting the possible generators of the monomial ideals occurring in a given factorization.
We conclude the paper by outlining, in Section 5, possible directions for future research into atomic factorizations in the monoids ${\mathcal{I}}(R)$ and ${\mathcal{M}{\rm on}}(R)$, with the present work serving as a foundation.
@@ 2.1. Factorization theory
Let $H$ be a commutative semigroup with an identity element, written multiplicatively. We denote by $H^{\times}$ the subgroup of invertible elements of $H$, also called units of $H$. If $H^{\times}=\{1\}$, we say that $H$ is reduced. An element $h\in H$ is cancellative if $h^{\prime},h^{\prime\prime}\in H$ and $hh^{\prime}=hh^{\prime\prime}$ implies $h^{\prime}=h^{\prime\prime}$, while it is unit-cancellative if $u\in H$ and $h=hu$ implies $u\in H^{\times}$. Moreover, we say that $H$ is cancellative (respectively, unit-cancellative) if all its elements are cancellative (respectively, unit-cancellative).
@@
Let $H$ be a cancellative monoid. An ideal of $H$ is a subset $J\subseteq H$ such that, for any $j\in J$, $jH\subseteq J$; they are also commonly referred to as $s$-ideals (semigroup ideals), especially when $H$ also carries a ring structure, to distinguish them from usual ring ideals. For ideals $J,K$ of $H$, denote by $(J:K)$ the ideal quotient $\{g\in\mathbf{q}(H):gK\subseteq J\}$. An ideal $J$ of $H$ is divisorial if $J=(H:(H:J))$. We say that $H$ is a Mori monoid if it satisfies the ascending chain condition on divisorial ideals. Furthermore, we denote by $\hat{H}:=\{g\in\mathbf{q}(H):hg^{n}\in H\text{\ for all\ }n\in\mathbb{N}\text{\ for some\ }h\in H\}$ the complete integral closure of $H$, and we say that $H$ is completely integrally closed if $H=\hat{H}$. Finally, we say that $H$ is a Krull monoid if it is Mori and completely integrally closed.
Let $H$ be a monoid, not necessarily cancellative. A submonoid $S\subseteq H$ is divisor-closed if, whenever $h\in S,h^{\prime}\in H$ and $h^{\prime}\mid h$, we have $h^{\prime}\in S$. For a subset $A\subseteq H$, we denote by $\llbracket A\rrbracket$ the smallest divisor-closed submonoid generated by $A$. A monoid homomorphism $\varphi:H\to H^{\prime}$, where $H^{\prime}$ is cancellative, is a divisor homomorphism if $h,h^{\prime}\in H$ and $\varphi(h)\mid\varphi(h^{\prime})$ in $H^{\prime}$ implies that $h\mid h^{\prime}$ in $H$.
An element $h\in H$ is prime if $h\notin H^{\times}$ and $h\mid h^{\prime}h^{\prime\prime}$ with $h^{\prime},h^{\prime\prime}\in H$ implies $h\mid h^{\prime}$ or $h\mid h^{\prime\prime}$. An element $h\in H$ is an atom if $h\notin H^{\times}$ and $h=h^{\prime}h^{\prime\prime}$ with $h^{\prime},h^{\prime\prime}\in H$ implies $h^{\prime}\in H^{\times}$ or $h^{\prime\prime}\in H^{\times}$; the set of all atoms of $H$ is denoted by $\mathscr{A}(H)$. A factorization of an element $a\in H$ is, up to units, a finite unordered sequence of atoms of $H$, say $u_{1},\dots,u_{n}$, such that $a=u_{1}\dots u_{n}$. The length of a factorization $z$ is the number of atoms occurring in it, and it is denoted by $\lvert z\rvert$. We denote by
@@ 2.2. Multivariate polynomial rings and their ideals
$R_{t}:=\{\sum_{m\in\mathbb{N}^{N}}r_{m}\mathbf{X}^{m}\colon r_{m}\in D\text{ and }m_{1}+\cdots+m_{N}=t\},$
$S_{t}:=\{\sum_{m\in\mathbb{N}^{N}}q_{m}\mathbf{X}^{m}\colon q_{m}\in K\text{ and }m_{1}+\cdots+m_{N}=t\}.$
@@
are graded rings, equipped with the standard grading. It is clear that $R_{t}\subseteq S_{t}$ for every $t$, that $R_{0}=D$ and $S_{0}=K$, and that $\deg X_{1}=\dots=\deg X_{N}=1$. It follows that every $f\in S$ can be uniquely written as $f=\sum_{t\geq 0}f_{t}$, with $f_{t}\in S_{t}$ for every $t$.
@@
Let ${\mathcal{I}}(R)$ be the semigroup of nonzero (ring) ideals of $R$, with the usual ideal multiplication. This semigroup is well known (see, e.g., [21, Prop. 2.1 (4)]) to be reduced and unit-cancellative, and thus a monoid in the sense specified in Section 2.1, being obviously commutative.
@@
It is of interest to understand how far is ${\mathcal{I}}(R)$ from being cancellative. We recall here a remarkable characterization of the cancellative elements of ${\mathcal{I}}(R)$ due to Anderson and Roitman, valid for any commutative ring $R$, not necessarily even a domain: an ideal $I$ of $R$ is cancellative if and only if it is a regular principal ideal at each localization at a maximal ideal [2]. If we consider $R=D[X_{1},\dots,X_{N}]$, with $D$ a noetherian domain and $N\geq 2$, it is easy to check that the cancellative elements of ${\mathcal{I}}(R)$ are precisely the principal ideals of $R$.
@@
be the set of the homogeneous components of degree $i$ of the elements of $I$, which is a $D$-module contained in $R_{i}$. For $I,J\in{\mathcal{I}}(R)$ and $i,j\in\mathbb{N}$,
@@
If $I=\langle g_{1},\dots,g_{n}\rangle$ is a nonzero finitely generated ideal of $R$ and $\operatorname{\mathsf{m-deg}}(I)=m\in\mathbb{N}$, then
@@
Throughout the rest of the paper, we will consider ideals of $R$ generated by polynomials in two indeterminates, say $X_{1}$ and $X_{2}$. To slightly ease the notation, we will rename these indeterminates to $X$ and $Y$ respectively. Note that if an ideal is generated only by polynomials in $X$ and $Y$, ideals occurring in any of its factorizations are necessarily also generated by polynomials in $X$ and $Y$ only. We will use this fact in the rest of the paper without further explanation.
Finally, let $\leq$ be the lexicographic order on monomials of $S$ defined by $X>Y>X_{3}>\dots>X_{N}$; we denote by $\operatorname{\mathsf{in}}(f)$ the initial monomial of $f\in S$ with respect to $\leq$, that is, the largest monomial occurring in $f$ with a non-zero coefficient.
@@ 2.3. A link between polynomial ideals and power monoids
Keeping the notation of the previous subsection, we let $R=D[X_{1},\dots,X_{N}]$, with $D$ an integral domain, be a polynomial ring in $N\geq 2$ indeterminates. As mentioned in the introduction, the study of ${\mathcal{I}}(R)$ from the perspective of factorization theory was initiated [21, Sect. 5], where it was shown that when ${\mathcal{I}}(R)$ is a $\mathsf{BF}$-monoid, it is fully elastic and each of its unions of sets of lengths equals $\mathbb{N}_{\geq 2}$ [21, Theorem 5.1].
@@ 3. Atoms and sets of lengths in ${\mathcal{I}}(R)$
Throughout this section, we maintain the notations introduced in Section 2.2, and assume that ${\mathcal{I}}(R)$ is a $\mathsf{BF}$-monoid. This assumption is clearly satisfied if, for instance, $D$ is noetherian (in particular, if $D=K$). Our aim is to find new classes of atoms in $\mathcal{I}(R)$, by extending the methods introduced in [21, Sect. 5] to further families of ideals of $R$, corresponding (via the map $\Phi$ in (2) from $\mathcal{P}_{\rm fin}(\mathbb{N})$ to ${\mathcal{I}}(R)$) to atoms of the reduced power monoid $\mathcal{P}_{{\rm fin},0}(\mathbb{N})$ of $\mathbb{N}$, i.e., the monoid of nonempty finite subsets of $\mathbb{N}$ containing 0 with setwise addition. Note that $\mathscr{A}(\mathcal{P}_{\rm fin}(\mathbb{N}))=\mathscr{A}(\mathcal{P}_{{\rm fin},0}(\mathbb{N}))\cup\{\{1\}\}$, since every finite set of positive integers $A=\{a_{1},\dots,a_{n}\}$ with $n\geq 2$ or $a_{1}\geq 2$ can be written as $A=(A-1)+\{1\}$. It is also worth noting that the monoid homomorphism $\Phi$ restricts to a monoid isomorphism between $\mathcal{P}_{{\rm fin},0}(\mathbb{N})$ and the ideals of $R$ generated, for some $m\in\mathbb{N}$, by $X^{m}$, $Y^{m}$ and a set of monomials in $X$ and $Y$ of degree $m$ (cf. [21, Example 5.13]).
The following lemma, generalizing the arguments used in [21], collects a number of necessary conditions for the existence of a nontrivial factorization for an ideal $I_{A}$, where $A\in\mathcal{P}_{{\rm fin},0}(\mathbb{N})$. They will be a key tool in the proofs of the main results of this section, namely, Theorem 3.3 and Theorem 3.7.
@@
Let $m\in\mathbb{N}^{+}$, $\mathcal{N}$ be a set of monomials of degree $m$ in $X$ and $Y$, and $I=\langle X^{m},Y^{m},\mathcal{N}\rangle$ be an ideal of $R$. Suppose that $I=\mathfrak{a}\cdot\mathfrak{b}$, for some $\mathfrak{a},\mathfrak{b}\in\mathcal{I}(R)$ such that $I\subsetneq\mathfrak{a}$, $\mathfrak{b}\subsetneq R$, $\operatorname{\mathsf{m-deg}}(\mathfrak{a})=d$, $\operatorname{\mathsf{m-deg}}(\mathfrak{b})=e$, ${\rm dim}_{K}(\mathfrak{a}_{K}[d])=r$ and ${\rm dim}_{K}(\mathfrak{b}_{K}[e])=s$. Then, the following hold.
@@
(v) Set $a_{1}=b_{1}=0$; then, for every $j\in\llbracket 1,r\rrbracket$ and every $k\in\llbracket 1,s\rrbracket$ we have $X^{m-(a_{j}+b_{k})}Y^{a_{j}+b_{k}}\in\{X^{m},Y^{m}\}\cup\mathcal{N}.$
@@
It follows from Remark 2.2 that $\operatorname{\mathsf{m-deg}}(I)=m$, and item (i) then follows from equation (1) and Lemma 2.5(i); item (ii) follows from [21, Lemma 5.8], while item (iv) is a consequence of item (iii). We are thus left to prove items (iii) and (v).
Since ${\rm dim}_{K}(\mathfrak{a}_{K}[d])=r$ and ${\rm dim}_{K}(\mathfrak{b}_{K}[e])=s$, we get from [21, Lemma 5.9] that there are bases $\{f_{1},\dots,f_{r}\}$ of $\mathfrak{a}_{K}[d]$ and $\{g_{1},\dots,g_{s}\}$ of $\mathfrak{b}_{K}[e]$ with $\operatorname{\mathsf{in}}(f_{1})>\operatorname{\mathsf{in}}(f_{2})>\dots>\operatorname{\mathsf{in}}(f_{r})$ and $\operatorname{\mathsf{in}}(g_{1})>\operatorname{\mathsf{in}}(g_{2})>\dots>\operatorname{\mathsf{in}}(g_{s})$. Note that the $\operatorname{\mathsf{in}}(f_{j})$ and $\operatorname{\mathsf{in}}(g_{k})$ are monomials of degree $d$ and $e$ respectively, for all $j\in\llbracket 1,r\rrbracket$ and $k\in\llbracket 1,s\rrbracket$. If $r=1$, we get from (i) that
@@
where we have used (iv) and (i). Since $f_{j}g_{k}\in I_{K}[m]={\rm Span}_{K}(\{X^{m},Y^{m}\}\cup\mathcal{N})$, we must also have $\operatorname{\mathsf{in}}(f_{j}g_{k})\in I_{K}[m]$, and the conclusion follows from the above equalities. ∎

### v2

@@ Abstract.
This research was funded in part by the Austrian Science Fund (FWF) [10.55776/DOC183 and 10.55776/ PAT9756623]. L. C. also acknowledges the European Union’s Horizon 2020 program (Marie Sklodowska-Curie grant 101021791) for the financial support on the early stages of the project. L. C. is a member of the National Group for Algebraic and Geometric Structures and their Applications (GNSAGA), a department of the Italian Mathematics Research Institute (INdAM)
@@ 1. Introduction
Our work contributes to the extensive literature on monoids of ideals: we refer to [29] and the monograph [17] for an overview, and to [33, 1, 30, 31] for more recent developments. A common approach in the literature is to focus on specific classes of ideals within a given ring, or semigroup ideals within a given monoid. Among these, divisorial ideals are perhaps the most prominent example. By suitably choosing a ring (or monoid) and a class of its ideals, one can obtain a fairly well-behaved monoid of ideals, and in some cases one is able to completely determine its structure. As an example, we recall two well-known results: a domain $R$ is Dedekind if and only if all its nonzero ideals are invertible, and this happens only if the monoid of all nonzero ideals is factorial, and a domain $R$ is Krull if and only if the monoid of its invertible divisorial ideals is factorial.
@@
In the paper [21], a first step was taken in studying the arithmetic properties of ${\mathcal{I}}(R)$, with $R$ a multivariate polynomial ring. In particular, the results collected in [21, Theorem 5.1] reveal a noticeable similarity between the arithmetic of ${\mathcal{I}}(R)$ and that of Krull monoids with infinite class group and prime divisors in each class, despite ${\mathcal{I}}(R)$ not even being transfer Krull. More specifically, ${\mathcal{I}}(R)$ is fully elastic and each of its unions of sets of lengths is the whole set $\mathbb{N}_{\geq 2}$.
Building on these initial results, we refine the demonstrative techniques and adapt them to broader contexts. Following a review of the necessary preliminaries in Section 2, we extend the methods introduced in [21] to find several new families of atoms in ${\mathcal{I}}(R)$. In particular, in Section 3.1 we highlight an interesting class built from sum-free subsets of $\mathbb{N}$, which includes some previously known examples.
Motivated by the problem of computing the length sets of certain ideals (see Theorem 3.7), in Section 4 we focus on the submonoid ${\mathcal{M}{\rm on}}(R)\subseteq{\mathcal{I}}(R)$ of monomial ideals of $R$ (i.e., ideals generated by monomials). We first provide some arithmetic results in the spirit of [21, Theorem 5.1], and then we determine several classes of atoms in this monoid, allowing us to compute the set of lengths of the families of ideals introduced in Theorem 3.7. The main technical ingredient that makes the study of factorizations in ${\mathcal{M}{\rm on}}(R)$ easier than in ${\mathcal{I}}(R)$ is Lemma 4.1 (see also Remark 4.2), greatly restricting the possible generators of the monomial ideals occurring in a given factorization.
We conclude the paper by outlining, in Section 5, possible directions for future research on atomic factorizations in the monoids ${\mathcal{I}}(R)$ and ${\mathcal{M}{\rm on}}(R)$, with the present work serving as a foundation.
@@ 2.1. Factorization theory
Let $H$ be a commutative semigroup with an identity element, written multiplicatively. We denote by $H^{\times}$ the subgroup of invertible elements of $H$, also called units of $H$. If $H^{\times}=\{1\}$, we say that $H$ is reduced. An element $a\in H$ is cancellative if $b,c\in H$ and $ab=ac$ implies $b=c$, while it is unit-cancellative if $b\in H$ and $a=ab$ implies $b\in H^{\times}$. Moreover, we say that $H$ is cancellative (respectively, unit-cancellative) if all its elements are cancellative (respectively, unit-cancellative).
@@
Let $H$ be a cancellative monoid. An ideal of $H$ is a subset $J\subseteq H$ such that, for any $a\in J$, $aH\subseteq J$; they are also commonly referred to as $s$-ideals (semigroup ideals), especially when $H$ also carries a ring structure, to distinguish them from usual ring ideals. For ideals $J,K$ of $H$, denote by $(J:K)$ the ideal quotient $\{x\in\mathbf{q}(H):xK\subseteq J\}$. An ideal $J$ of $H$ is divisorial if $J=(H:(H:J))$. We say that $H$ is a Mori monoid if it satisfies the ascending chain condition on divisorial ideals. Furthermore, we denote by $\hat{H}:=\{x\in\mathbf{q}(H):ax^{n}\in H\text{\ for all\ }n\in\mathbb{N}\text{\ for some\ }a\in H\}$ the complete integral closure of $H$, and we say that $H$ is completely integrally closed if $H=\hat{H}$. Finally, we say that $H$ is a Krull monoid if it is Mori and completely integrally closed.
Let $H$ be a monoid, not necessarily cancellative. A submonoid $S\subseteq H$ is divisor-closed if, whenever $a\in S,b\in H$ and $b\mid h$, we have $b\in S$. For a subset $A\subseteq H$, we denote by $\llbracket A\rrbracket$ the smallest divisor-closed submonoid generated by $A$. A monoid homomorphism $\varphi:H\to H^{\prime}$, where $H^{\prime}$ is cancellative, is a divisor homomorphism if $a,b\in H$ and $\varphi(a)\mid\varphi(b)$ in $H^{\prime}$ implies that $a\mid b$ in $H$.
An element $a\in H$ is prime if $a\notin H^{\times}$ and $a\mid bc$ with $b,c\in H$ implies $a\mid b$ or $a\mid c$. An element $a\in H$ is an atom if $a\notin H^{\times}$ and $a=bc$ with $b,c\in H$ implies $b\in H^{\times}$ or $c\in H^{\times}$; the set of all atoms of $H$ is denoted by $\mathscr{A}(H)$. A factorization of an element $a\in H$ is, up to units, a finite unordered sequence of atoms of $H$, say $u_{1},\dots,u_{n}$, such that $a=u_{1}\dots u_{n}$. The length of a factorization $z$ is the number of atoms occurring in it, and is denoted by $\lvert z\rvert$. We denote by
@@
  • $\mathcal{U}_{k}(H):=\bigcup_{k\in L\in\mathcal{L}(H)}L$ the union of sets of lengths of $H$ containing $k\in\mathbb{N}$;
@@ 2.2. Multivariate polynomial rings and their ideals
$R_{t}:=\Biggl\{\sum_{m\in\mathbb{N}^{N}}r_{m}\mathbf{X}^{m}\colon r_{m}\in D\text{ and }m_{1}+\cdots+m_{N}=t\Biggr\},$
$S_{t}:=\Biggl\{\sum_{m\in\mathbb{N}^{N}}q_{m}\mathbf{X}^{m}\colon q_{m}\in K\text{ and }m_{1}+\cdots+m_{N}=t\Biggr\}.$
@@
are graded rings, equipped with the standard grading, so that $\deg X_{1}=\dots=\deg X_{N}=1$. It is clear that $R_{t}\subseteq S_{t}$ for every $t$, that $R_{0}=D$ and $S_{0}=K$, and that every $f\in S$ can be written uniquely as $f=\sum_{t\geq 0}f_{t}$, with $f_{t}\in S_{t}$ for every $t$.
@@
Let ${\mathcal{I}}(R)$ be the semigroup of nonzero (ring) ideals of $R$, with the usual ideal multiplication. This semigroup is commutative, has an identity element ($R$ itself) and it is reduced. Moreover, if $R$ satisfies the Krull Intersection Theorem, i.e., if $\bigcap_{n\geq 0}I^{n}=\{0\}$ for every $I\in{\mathcal{I}}(R)\setminus\{R\}$ (this holds, for example, if $D$ is noetherian), then ${\mathcal{I}}(R)$ is also unit-cancellative by [21, Prop. 2.1(4)], and thus a monoid in the sense specified in Section 2.1.
@@
It is of interest to understand how far ${\mathcal{I}}(R)$ is from being cancellative. We recall here a remarkable characterization of the cancellative elements of ${\mathcal{I}}(R)$ due to Anderson and Roitman, valid for any commutative ring $R$, not necessarily even a domain: an ideal $I$ of $R$ is cancellative if and only if it is a regular principal ideal at each localization at a maximal ideal [2]. If we consider $R=D[X_{1},\dots,X_{N}]$, with $D$ a noetherian domain and $N\geq 2$, it is easy to check that the cancellative elements of ${\mathcal{I}}(R)$ are precisely the principal ideals of $R$.
@@
be the set of homogeneous components of degree $i$ of the elements of $I$, which is a $D$-module contained in $R_{i}$. For $I,J\in{\mathcal{I}}(R)$ and $i,j\in\mathbb{N}$,
@@
If $I=\langle g_{1},\dots,g_{n}\rangle$ is a finitely generated nonzero ideal of $R$ and $\operatorname{\mathsf{m-deg}}(I)=m\in\mathbb{N}$, then
@@
Throughout the rest of the paper, we will consider ideals of $R$ generated by polynomials in two indeterminates, say $X_{1}$ and $X_{2}$. To slightly ease the notation, we will rename these indeterminates to $X$ and $Y$ respectively.
Finally, let $\leq$ be the lexicographic order on monomials of $S$ defined by $X>Y>X_{3}>\dots>X_{N}>1$. We denote by $\operatorname{\mathsf{in}}(f)$ the initial monomial of a nonzero $f\in S$ with respect to $\leq$, that is, the largest monomial with respect to the lexicographic order occurring in $f$ with a nonzero coefficient. Following [13], we set $\operatorname{\mathsf{in}}(0)=0$.
@@
In the notation of Lemma 2.5, if we further assume $(IJ)_{K}[m]\subseteq K[X,Y]$, then for every $f\in I_{K}[d]$ and every $g\in J_{K}[e]$, we have $\operatorname{\mathsf{in}}(f),\,\operatorname{\mathsf{in}}(g)\in K[X,Y]$. In fact, if there exists a nonzero $f\in I_{K}[d]$ such that $X_{i}$ divides $\operatorname{\mathsf{in}}(f)$ for some $i\in\llbracket 3,N\rrbracket$, then for any nonzero $g\in J_{K}[e]$ (one always exists as $J$ is a nonzero ideal) we have $\operatorname{\mathsf{in}}(fg)=\operatorname{\mathsf{in}}(f)\operatorname{\mathsf{in}}(g)\not\in K[X,Y]$, and so $fg\in(IJ)_{K}[m]\setminus K[X,Y]$, contradicting the assumption.
@@ 2.3. A link between polynomial ideals and power monoids
Keeping the notation of the previous subsection, we let $R=D[X_{1},\dots,X_{N}]$, with $D$ an integral domain, be a polynomial ring in $N\geq 2$ indeterminates. As mentioned in the introduction, the study of ${\mathcal{I}}(R)$ from the perspective of factorization theory was initiated in [21, Sect. 5], where it was shown that when ${\mathcal{I}}(R)$ is a $\mathsf{BF}$-monoid, it is fully elastic and each of its unions of sets of lengths equals $\mathbb{N}_{\geq 2}$ [21, Theorem 5.1].
@@ 3. Atoms and sets of lengths in ${\mathcal{I}}(R)$
Throughout this section, we maintain the notation introduced in Section 2.2, and assume that ${\mathcal{I}}(R)$ is a $\mathsf{BF}$-monoid. This assumption is satisfied if, for instance, $D$ is noetherian (in particular, if $D=K$). Our aim is to find new classes of atoms in ${\mathcal{I}}(R)$, by extending the methods introduced in [21, Sect. 5] to further families of ideals of $R$, corresponding (via the map $\Phi$ in (2) from $\mathcal{P}_{\rm fin}(\mathbb{N})$ to ${\mathcal{I}}(R)$) to atoms of the reduced power monoid $\mathcal{P}_{{\rm fin},0}(\mathbb{N})$ of $\mathbb{N}$, i.e., the monoid of nonempty finite subsets of $\mathbb{N}$ containing 0 with setwise addition. Note that $\mathscr{A}(\mathcal{P}_{\rm fin}(\mathbb{N}))=\mathscr{A}(\mathcal{P}_{{\rm fin},0}(\mathbb{N}))\cup\{\{1\}\}$, since every finite set of positive integers $A=\{a_{1},\dots,a_{n}\}$ with $n\geq 2$ or $a_{1}\geq 2$ can be written as $A=(A-1)+\{1\}$. It is also worth noting that the monoid homomorphism $\Phi$ restricts to a monoid isomorphism between $\mathcal{P}_{{\rm fin},0}(\mathbb{N})$ and the ideals of $R$ generated, for some $m\in\mathbb{N}$, by $X^{m}$, $Y^{m}$ and a set of monomials in $X$ and $Y$ of degree $m$ (cf. [21, Example 5.13]).
The following lemma, generalizing the arguments used in [21], collects a number of necessary conditions for the existence of a nontrivial factorization for an ideal $I_{A}$, where $A\in\mathcal{P}_{{\rm fin},0}(\mathbb{N})$. They will be a key tool in the proofs of the main results of this section, namely Theorem 3.3 and Theorem 3.7.
@@
Let $m\in\mathbb{N}^{+}$, $\mathcal{G}$ be a set of monomials of degree $m$ in $X$ and $Y$, and let $I=\langle X^{m},Y^{m},\mathcal{G}\rangle$ be an ideal of $R$. Suppose that $I=\mathfrak{a}\cdot\mathfrak{b}$, for some $\mathfrak{a},\mathfrak{b}\in{\mathcal{I}}(R)$ such that $I\subsetneq\mathfrak{a}$, $\mathfrak{b}\subsetneq R$, $\operatorname{\mathsf{m-deg}}(\mathfrak{a})=d$, $\operatorname{\mathsf{m-deg}}(\mathfrak{b})=e$, ${\rm dim}_{K}(\mathfrak{a}_{K}[d])=r$ and ${\rm dim}_{K}(\mathfrak{b}_{K}[e])=s$. Then, the following hold.
@@
(v) Set $a_{1}=b_{1}=0$; then, for every $j\in\llbracket 1,r\rrbracket$ and every $k\in\llbracket 1,s\rrbracket$ we have $X^{m-(a_{j}+b_{k})}Y^{a_{j}+b_{k}}\in\{X^{m},Y^{m}\}\cup\mathcal{G}.$
@@
It follows from Remark 2.2 that $\operatorname{\mathsf{m-deg}}(I)=m$, and item (i) then follows from equation (1) and Lemma 2.5(i); item (ii) follows from [21, Lemma 5.8], while item (iv) is a consequence of (iii) and 2.6.
Since ${\rm dim}_{K}(\mathfrak{a}_{K}[d])=r$ and ${\rm dim}_{K}(\mathfrak{b}_{K}[e])=s$, we get from [21, Lemma 5.9] that there are bases $\{f_{1},\dots,f_{r}\}$ of $\mathfrak{a}_{K}[d]$ and $\{g_{1},\dots,g_{s}\}$ of $\mathfrak{b}_{K}[e]$ with $\operatorname{\mathsf{in}}(f_{1})>\operatorname{\mathsf{in}}(f_{2})>\dots>\operatorname{\mathsf{in}}(f_{r})$ and $\operatorname{\mathsf{in}}(g_{1})>\operatorname{\mathsf{in}}(g_{2})>\dots>\operatorname{\mathsf{in}}(g_{s})$. Note that $\operatorname{\mathsf{in}}(f_{j})$ and $\operatorname{\mathsf{in}}(g_{k})$ are monomials of degree $d$ and $e$ respectively, for all $j\in\llbracket 1,r\rrbracket$ and $k\in\llbracket 1,s\rrbracket$. If $r=1$, we get from (i) that
@@
where we have used (iv) and (i). Since $f_{j}g_{k}\in I_{K}[m]={\rm Span}_{K}(\{X^{m},Y^{m}\}\cup\mathcal{G})$, we must also have $\operatorname{\mathsf{in}}(f_{j}g_{k})\in I_{K}[m]$, and the conclusion follows from the above equalities. ∎